Solve the following trigonometric equations. · cos(x) = } · cos(x) + V3 sin(x) = V2
Answer #1

cold (n) = 4 - or cos(x) = cos(na I 7 ) where nez 1. - x= nath &NER leachner I cos(nat - cos(n) + VB sin(n) = 12 or, cos(n) +